# xLSTM-Mixer: Multivariate Time Series Forecasting by Mixing via Scalar Memories

### v2

@@ 3.3 Key Component 3: Multi-View Mixing
To further regularize the training of the sLSTM as with the linear projections, we compute forecasts from the original embedding $\bm{x}^{\text{up}}$ as well as the reversed embedding $\bm{\widehat{x}}^{\text{up}}$, where the order of the latent dimensions including the representation of $\bm{\eta}$ is inverted. Learning forecasts $\bm{y}^{\prime}$ and $\bm{y}^{\prime\prime}$ for both views while sharing weights helps learn better representations. Such multi-task learning settings are known to benefit training (Zhang & Yang, 2022). The final forecast is obtained by a linear projection $\operatorname{FC}^{\text{view}}$ of the two concatenated forecasts, again per-variate. Specifically, we compute:
[…]
The final forecast is obtained after de-normalizing the reconciled forecasts as $\bm{y}=\operatorname{RevIN}^{-1}(\bm{y}^{\text{norm}})$.
4 Experimental Evaluation
@@
Our intention here is to evaluate the forecasting capabilities of xLSTM-Mixer, aiming to provide comprehensive insights into its performance. To this end, we conducted a series of experiments with the primary focus on long-term forecasting, following the work of Das et al. (2023) and Chen et al. (2023c). An evaluation of xLSTM-Mixer’s competitiveness in short-term forecasting on the PEMS dataset is provided in Sec. A.2. Additionally, we perform an extensive model analysis consisting of an ablation study to identify the contributions of individual components of xLSTM-Mixer, followed by an inspection of the initial embedding tokens, a hyperparameter sensitivity analysis, and an investigation into its robustness.
Dataset Source Domain Horizons Sampling #Variates
Weather Zhou et al. (2021) Weather 96–720 10 min 21
Electricity Zhou et al. (2021) Power Usage 96–720 1 hour 321
Traffic Wu et al. (2021) Traffic Load 96–720 1 hour 862
ETT Zhou et al. (2021) Power Production 96–720 15&60 min 7
Datasets.   We generally follow the established benchmark procedure of Wu et al. (2021) and Zhou et al. (2021) for best backward and future comparability. The datasets we thus used are summarized in Tab. 1.
Training.   We follow standard practice in the forecasting literature by evaluating long-term forecasts using the mean squared error (MSE) and the mean absolute error (MAE). Based on our experiments, we used the MAE as the training loss function since it yielded the best results. The datasets were standardized for consistency across features. Further details on hyperparameter selection, metrics, and the implementation can be found in Sec. A.1.
Baseline Models.   We compare xLSTM-Mixer to the recurrent models xLSTMTime (Alharthi & Mahmood, 2024) and LSTM (Hochreiter & Schmidhuber, 1997); multi-perceptron (MLP) based models TimeMixer (Wang et al., 2024a), TSMixer (Chen et al., 2023c), DLinear (Zeng et al., 2023), and TiDE (Das et al., 2023); the Transformers PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2023), FEDFormer (Zhou et al., 2022), and Autoformer (Wu et al., 2021); and the convolutional architectures MICN (Wang et al., 2022) and TimesNet (Wu et al., 2022).
@@ 4.1 Long-Term Time Series Forecasting
Table 2: xLSTM-Mixer is effective in long-term forecasting. The results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined. Wins for each model out of all 28 settings are shown at the bottom.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
@@
iTransformer
FEDFormer
Autoformer
MICN
TimesNet
1997a
@@
0
0
0
0
0
0
0
0
  • a
    Taken from Wu et al. (2022).
@@
Figure 2: Example Forecasts Across Models and Datasets. This figure shows example forecasts on the Weather and ETTm1 datasets for multiple models. The blue lines represent the ground truth target, and the orange lines show the predictions. The first panel illustrates the forecast from xLSTM-Mixer, while the second panel shows the forecast extracted before the up-projection step, highlighting the effectiveness of our added components. Comparisons with xLSTMTime, TimeMixer, and PatchTST confirm the favorable performance of xLSTM-Mixer relative to these baseline models.
@@
We present the performance of xLSTM-Mixer compared to prior models in Tab. 2. As shown, xLSTM-Mixer consistently delivers highly accurate forecasts across a wide range of datasets. It achieves the best results in 18 out of 28 cases for MSE and 22 out of 28 cases for MAE, demonstrating its superior performance in long-term forecasting. In particular, xLSTM-Mixer exhibits exceptional forecasting accuracy, as evidenced particularly by its strong MAE performance across all datasets. Notably, on Weather, xLSTM-Mixer reduces the MAE by 2% compared to xLSTMTime and 4.6% compared to TimeMixer. Similarly, for ETTm1, xLSTM-Mixer outperforms TimeMixer by 2.4% in MAE and shows a strong competitive edge over xLSTMTime. Although xLSTM-Mixer performs slightly less well on the Traffic and ETTh2 datasets, where it encounters challenges with handling outliers, it remains highly competitive and outperforms the majority of baseline models. This suggests that despite these few cases, xLSTM-Mixer can consistently deliver state-of-the-art performance in long-term forecasting. A qualitative inspection of several baseline models, including the initial forecast extracted before the sLSTM refinement, is shown in Fig. 2. In this comparison, the lookback window and forecasting horizon are both fixed at 96.
@@ 4.2 Model Analysis
Ablation Study.
To assess the contributions of each component in xLSTM-Mixer to its strong overall forecast performance, we conducted an extensive ablation study with the results listed in Tab. 3. Each configuration represents a different combination of the four key components: mixing time with NLinear, using sLSTM blocks, learning an initial embedding token, and multi-view mixing. We evaluated the performance using the MSE and MAE across the prediction lengths $\{96,192,336,720\}$.
Table 3: Ablation Study on the Weather and ETTm1 Datasets. The MSE (↓) and MAE (↓) are reported for each numbered configuration across all four forecast horizons. The best results are highlighted in bold red, while the second-best results are blue and underlined.
@@
Mix Time
@@
The full version of xLSTM-Mixer (#1), which integrates all components, achieves the best performance overall. However, we also observe that some configurations of xLSTM-Mixer, which exclude specific components, remain competitive. For instance, #3, which excludes the initial embedding token, still performs reasonably well. This suggests that while it contributes positively to the overall performance, the model can sometimes still achieve competitive results without it. In general, removing any specific component leads to a performance drop. For example, removing the time mixing (#7) increases the MAE by 3.4% on ETTm1 at length 96 or 2.8% at length 192, highlighting its critical role in capturing intratemporal dependencies. When we now omit everything except for time mixing on Weather at 192, we suffer a 13.7% performance decrease. In summary, the ablation study confirms that all components of xLSTM-Mixer contribute to its effectiveness, with the full configuration yielding the best results. Furthermore, we identified the sLSTM blocks and time-mixing as critical components for ensuring high accuracy across datasets and prediction lengths.
Figure 3: Initial tokens capture dataset characteristics. The plot illustrates the learned tokens across multiple datasets and prediction lengths. The lookback length is set to 96 for all evaluations. For clarity and the high noise levels in ETTm1 and ETTh2, we only show three seeds for Weather.
Initial Token Embedding.
We qualitatively inspect decodings of the initial embedding tokens $\bm{\eta}$ on multiple datasets to further understand and interpret the initializations learned by xLSTM-Mixer. $\bm{\eta}$ are decoded to a forecast $\bm{y}$ by transforming them through the sLSTM stack $\mathcal{S}$ and applying multi-view mixing. The resulting output of $\operatorname{FC}^{\text{view}}$ can then be interpreted as the conditioning forecast used to initialize the sLSTM blocks. Fig. 3 shows the dataset-specific patterns the initial embedding tokens have learned on Weather, ETTm1, and ETTh2 for various prediction horizons. With increasing prediction horizons, we observe longer spans of time, eventually revealing underlying seasonal patterns and respective dataset dynamics.
Figure 4: Sensitivity analysis of the xLSTM hidden dimension $D$. The shaded area indicates one standard deviation from the mean MSE (↓). Increasing the up-projection dimension becomes beneficial with increasing prediction length.
Sensitivity to xLSTM Hidden
Dimension.
In Fig. 4, we visualize the performance of xLSTM-Mixer on the Electricity dataset with increasing sLSTM embedding (hidden) dimension realized by $\operatorname{FC}^{\text{up}}$. The results indicate that larger hidden dimensions consistently enhance the model’s performance, particularly for longer forecast horizons. This suggests that a larger embedding dimension enables xLSTM-Mixer to capture better the higher complexity of the time series data over extended horizons, leading to improved forecasting accuracy.
Robustness to Lookback Length.
Fig. 5 illustrates the performance of xLSTM-Mixer across varying lookback lengths and prediction horizons. We observe that xLSTM-Mixer can effectively utilize longer lookback windows than the baselines, especially when compared to transformer-based models. This advantage stems from xLSTM-Mixer’s avoidance of self-attention, allowing it to handle extended lookback lengths efficiently. Additionally, xLSTM-Mixer demonstrates stable and consistent performance with low variance. These results confirm that increasing the lookback length improves forecasting accuracy and enhances robustness, particularly for longer prediction horizons.
Figure 5: Increasing the lookback window increases forecasting performance, with xLSTM-Mixer virtually always providing the best results. Shows the mean MSE (↓) and std. dev. on ETTm1.

### v3

@@ 3.3 Multi-View Mixing
To further regularize the training of the sLSTM as with the linear projections, we compute forecasts from the original embedding $\bm{x}^{\text{up}}$ as well as the reversed embedding $\bm{\widehat{x}}^{\text{up}}$, where the order of the latent dimensions including the representation of $\bm{\eta}$ is inverted. Learning forecasts $\bm{y}^{\prime},\bm{y}^{\prime\prime}\in\mathbb{R}^{V\times D}$ for both views while sharing weights helps learn better representations. Such multi-task learning settings are known to benefit training (Zhang and Yang, 2022). It can also be viewed as ensembling over different variate orderings with weight sharing. The final forecast is obtained by a linear projection $\operatorname{FC}^{\text{view}}:\mathbb{R}^{D}\times\mathbb{R}^{D}\to\mathbb{R}^{H}$ of the two forecasts, again per-variate. Specifically, we compute $\bm{y}^{\text{norm}}=\operatorname{FC}^{\text{view}}\left(\bm{y}^{\prime},\bm{y}^{\prime\prime}\right)$, where $\bm{y}^{\prime}=\mathcal{S}(\bm{x}^{\text{up}})$ and $\bm{y}^{\prime\prime}=\mathcal{S}(\bm{\widehat{x}}^{\text{up}})$. The final forecast is obtained after de-normalizing the reconciled forecasts as $\bm{y}=\operatorname{RevIN}^{-1}(\bm{y}^{\text{norm}})$.
4 Experiments
@@
We conduct a series of experiments to evaluate the forecasting capabilities of xLSTM-Mixer, aiming to provide comprehensive insights into its performance. Our primary focus is on long-term forecasting, following the works of Das et al. (2023); Chen et al. (2023c); Lin et al. (2024), and Liu et al. (2025). Further tasks are explored in Sec. 4.3. Additionally, we perform an extensive model analysis, including visualizations of the initial embedding tokens, hyperparameter sensitivity, and performance measurement. Finally, an ablation study identifies the contributions of the individual components.
Datasets.   We generally follow the established benchmark procedure of Wu et al. (2021) and Zhou et al. (2021) for best backward and future comparability. The datasets we thus used are provided as an overview in App. D. Training.   We follow standard practice in the forecasting literature by evaluating long-term forecasts using mean squared error (MSE) and mean absolute error (MAE). Based on our experiments, we used MAE as the training loss function since it yielded the best results. The datasets were standardized for consistency across features. In addition, we conducted all experiments three times and reported the averaged values. Further details on hyperparameter selection, metrics, and implementation can be found in App. C. Baseline Models.   We compare xLSTM-Mixer to the recurrent models xLSTMTime (Alharthi and Mahmood, 2024) and LSTM (Hochreiter and Schmidhuber, 1997); mixer models TimeMixer++ (Wang et al., 2025a), TimeMixer (Wang et al., 2024a), and TSMixer (Chen et al., 2023c); MLP-based models CycleNet/MLP (Lin et al., 2024), DLinear (Zeng et al., 2023), and TiDE (Das et al., 2023); the SSMs S-Mamba (Wang et al., 2025b) and Chimera (Behrouz et al., 2024); the Transformers PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023); the convolutional architectures ModernTCN (Donghao and Xue, 2023) and TimesNet (Wu et al., 2022a); and the pretrained zero-shot forecasters Timer-XL (Liu et al., 2025) and MoiraiBase (Woo et al., 2024). On choosing lookback lengths $L$.   Some prior works on long-term forecasting fix the lookback window $L$ for the sake of a fair comparison. However, Abdelmalak et al. (2025) and Brigato et al. (2025) argue that fixing especially to 96, which is common in today’s benchmarks, can substantially distort the comparisons. When allowed to tune each model’s input length, most baselines improve and simple linear or MLP backbones close much of the gap to Transformers, whereas at fixed $L=96$ they appear artificially weak. We agree with these findings and therefore provide results for optimal hyperparameters for all baselines and xLSTM-Mixer in our experiments. See also 4(b) in Sec. 4.4 for a sensitivity analysis on $L$.
@@ 4.1 Long-Term Time Series Forecasting
Table 1: xLSTM-Mixer is effective in long-term forecasting. The results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted bold red and second-best blue and underlined. Wins for each model out of all 28 settings are shown at the bottom. Full results are provided in App. E.
Models Recurrent Mixer MLP State Space Trans. Conv. Pretrained*
xLSTM-
@@
Mod.TCN
Timer-XL
MoiraiBase
Mixer
1997 $\dagger$
@@
    Zero-shot forecasting.
  • $\dagger$
    Taken from Wu et al. (2022a).
  • $\ddagger$
    Traffic/PEMS are often used during pre-training (Liu et al., 2025). Thus, no zero-shot results are available.
@@
Figure 3: xLSTM-Mixer provides convincing forecasts. This figure shows example forecasts on the Weather and ETTm1 datasets for multiple models with lookback windows and forecasting horizons fixed at 96. The first panel illustrates the forecast from xLSTM-Mixer, while the second shows the initial forecast extracted before the up-projection step, highlighting the effectiveness of our added components. Comparisons with further baselines are provided for context.
@@
We present the performance of xLSTM-Mixer compared to prior models in Tab. 1. The full results and standard deviations are provided in App. E. As shown, xLSTM-Mixer consistently delivers highly accurate forecasts across a wide range of datasets. It achieves the best results in 11 out of 28 cases for MSE and 16 out of 28 cases for MAE, demonstrating its superior performance in long-term forecasting. xLSTM-Mixer defines a new state-of-the-art on six out of seven datasets. This shows that xLSTM-Mixer consistently delivers excellent performance in long-term forecasting. A qualitative comparison with several baselines, including the initial forecast extracted before the sLSTM refinement, is shown in Fig. 3. Here, both the lookback window and the forecasting horizon are fixed at 96.
Significance Testing.   In addition to the common practice of evaluating by counting wins across models and datasets, we perform statistical testing following the well-known practice of Demšar (2006). First, a Friedman test ensures the model’s performances follow different distributions ($p<10^{-10}$). We can then perform a Conover post-hoc test while adjusting for the family-wise error rate using Holm’s method. We must restrict this comparison to one metric (MSE) and horizon (96 steps), to avoid the strongly correlated results for different horizons to artificially inflate significance levels. At a significance threshold of $p=0.05$, we find that xLSTM-Mixer is statistically significantly better than all other methods, except for xLSTMTime. Yet, the difference in average rank is still impressive, with 1.5 for xLSTM-Mixer and 4.0 for xLSTMTime. Furthermore, xLSTMTime is statistically inseparable from many other models, namely TimeMixer, TSMixer, ModernTCN, Chimera, PatchTST, and TiDE. See Fig. 5 in App. E for a visual critical difference (CD) diagram.
4.2 Forecasting on the GIFT-Eval Benchmark
To showcase xLSTM-Mixer’s versatility beyond multivariate long-term point forecasting, we evaluate it on GIFT-Eval (Aksu et al., 2024). It is a comprehensive benchmark for general time-series forecasting spanning diverse domains, frequencies, horizons, and both univariate and multivariate regimes. The benchmark emphasizes probabilistic assessment and supports standardized (including zero-shot) evaluation protocols. Following the official setup, we report aggregated Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS), where lower is better. To extend xLSTM-Mixer to probabilistic forecasting capabilities, we equip it with a head predicting quantiles, yielding full predictive distributions rather than point estimates.
Results are summarized in Tab. 2 and detailed in App. F. By CRPS, xLSTM-Mixer is the top purely supervised model (i.e., without using the GIFT-Eval pretraining corpus) and ranks 2nd overall. This demonstrates that although xLSTM-Mixer is designed for multivariate forecasting, it remains highly competitive among methods across univariate and multivariate settings and for both short and long horizons in the probabilistic regime.
Table 2: xLSTM-Mixer excels on GIFT-Eval. The table reports the top 10 across all categories.
Model MASE $\downarrow$ CRPS $\downarrow$ Rank (CRPS) $\downarrow$
TiRex $0.724$ $0.498$ 11
xLSTM-Mixer (ours) $0.780$ $0.510$ 22
TEMPO_ensemble $0.862$ $0.514$ 33
Toto_Open_Base_1.0 $0.750$ $0.517$ 44
TabPFN-TS $0.771$ $0.544$ 55
YingLong_300m $0.798$ $0.548$ 66
timesfm_2_0_500m $0.758$ $0.550$ 77
YingLong_110m $0.809$ $0.557$ 88
sundial_base_128m $0.750$ $0.559$ 99
YingLong_50m $0.822$ $0.567$ $10$
Significance Testing.   As in the previous Sec. 4.1, we rigorously test for the significance of the improvement of xLSTM-Mixer over the existing methods. After the initial Firedman test on the MSE, we see that xLSTM-Mixer is significantly better than all other 38 methods, except for sundial_base_128m, TiRex, and TTM-R2-Finetuned. See Fig. 6 in App. F for a CD diagram.
4.3 Outlook on Classification
To illustrate performance beyond forecasting, we evaluate xLSTM-Mixer on time-series classification by using it as an embedding model. For this, we replace the final projection layer with a single classification head. xLSTM-Mixer achieves very strong performance on standard time-series classification benchmarks (cf. App. J). These findings suggest that approaches reconciling recurrent and mixing architectures, such as xLSTM-Mixer, are highly flexible yet powerful.
@@ 4.4 Model Analysis
Initial Token Embeddings.   We qualitatively inspect decodings of the initial embedding tokens $\bm{\eta}$ on multiple datasets to further understand and interpret the initializations learned by xLSTM-Mixer. $\bm{\eta}$ are decoded to a forecast $\bm{y}$ by transforming them through the sLSTM stack $\mathcal{S}$ and applying multi-view mixing. The resulting output of $\operatorname{FC}^{\text{view}}$ can then be interpreted as the conditioning time series used to initialize the sLSTM blocks. Fig. 7 in Appendix App. G shows the dataset-specific patterns the initial embedding tokens have learned for various horizons $H$. Increasing prediction horizons eventually reveal underlying seasonal patterns and respective dataset dynamics.
Model Efficiency.   To assess the computational resources required for using xLSTM-Mixer, we measured the average wall-clock time and peak graphics card memory required to perform a training step. Fig. 1 shows how this changes over multiple lookback lengths $T$ and two datasets at a forecast horizon of $H=336$. Compared to baselines, xLSTM-Mixer scales extremely favorably in $T$, only exhibiting a negligible increase in time and memory requirements. While computations take slightly longer for larger lookback sizes, the increase is much smaller than for Transformer-based models. One advantage of TimeMixer was its efficiency over Transformers, upon which xLSTM-Mixer now significantly improves by requiring one to two orders of magnitude less memory.
Sensitivity to the xLSTM Hidden Dimension.   In 4(a), we visualize the performance of xLSTM-Mixer on the Electricity dataset with increasing sLSTM embedding hidden dimension $D$ realized by $\operatorname{FC}^{\text{up}}$. The results indicate that a larger $D$ enables xLSTM-Mixer to better capture the complexity of the series over extended horizons, leading to improved forecasting accuracy.
(a) Increasing the hidden dimension $D$ becomes increasingly beneficial at longer prediction lengths $H$. Shown is the MSE for varying $D$ on the Electricity dataset.
(b) Increasing the lookback window $T$ increases forecasting performance, with xLSTM-Mixer virtually always providing the best results. Shows the MSE (↓) on the ETTm1 dataset.
Figure 4: Impact of model parameters on forecasting performance.
Robustness to the Lookback Length.   4(b) illustrates the performance of xLSTM-Mixer across varying lookback lengths $T$ and prediction horizons $H$. Note that we had to rerun some experiments for TimeMixer at $T=720$ with varying seeds since many training runs diverged. We observe that xLSTM-Mixer can effectively utilize longer lookback windows than the baselines, especially when compared to transformer-based models. This advantage stems from xLSTM-Mixer’s avoidance of self-attention, allowing it to handle extended lookback lengths efficiently. On short prediction lengths with $T\in\{96,192\}$, information of more than 768 time steps in the past becomes redundant to inform the comparatively short forecast, causing models to deteriorate slightly. On longer horizons, increasingly farther lookbacks become useful for forecasting. Additionally, xLSTM-Mixer demonstrates stable and consistent performance with low variance across scales.
xLSTM-Mixer Captures Cross-Variate Patterns.   To determine if our chosen ordering of variates is viable, we conduct additional experiments investigating the impact of different variate permutations in Tab. 12 in App. H. While we observe that the performance does depend to a certain extent on the ordering of variates, this does not pose a significant limitation in practice since the standard ordering provided with each dataset already yields strong forecasts. Moreover, our attribution analysis (cf. Fig. 8) confirms that xLSTM-Mixer effectively captures cross-variate interactions.
Ensembling over More than Two Views.   Following the perspective that multi-view mixing is an ensemble over variate orderings, we extend our method to accommodate additional randomly permuted views beyond the original and reverse order. To this end, $\operatorname{FC}^{\text{view}}$ is extended accordingly. The full results in App. I show that this does not improve the modeling accuracy of xLSTM-Mixer.
4.5 Ablation Studies
Table 3: Each component of xLSTM-Mixer is essential for its overall strong performance. The notation follows Tab. 1. Results are averages over three seeds.
@@ 4.5 Ablation Studies
#11
#12
#13
#14
Time Mixing
DLinear
xLSTM type
@@
mLSTM
LSTM
GRU
sLSTM
sLSTM
sLSTM
sLSTM
sLSTM
None
sLSTM
sLSTM
sLSTM
sLSTM
Recurr. order
Variates
Variates
Variates
@@
To assess the contributions of each component in xLSTM-Mixer to its strong forecast performance, we conducted ablation studies of thirteen different model configurations with the results listed in Tab. 3. Each configuration represents a different combination of the four key components: mixing time with NLinear/DLinear, using sLSTM/mLSTM/LSTM/GRU for joint mixing, striding over the variate or time dimension, learning initial embedding tokens $\bm{\eta}$, and multi-view mixing.
@@
The full version of xLSTM-Mixer (#1), which integrates all components, achieves the best performance overall. Ablating components of xLSTM-Mixer each causes both error metrics to increase, entailing that they contribute positively to the overall architecture. Specifically, omitting xLSTM (only having LSTM, #3) raises the MAE by 6.2%/the MSE by 7.0%, variate recurrence (#5) by 4.3%/4.7%, learning initial token embedding (#7) by 0.4%/0.7%, view mixing (#8) by 0.6%/0.7%, and time mixing (#11) by 2.7%/3.1%. For example, removing the time mixing (#11) increases the MAE by 3.4% on ETTm1 at length 96 or 3.1% at length 192, highlighting its critical role in capturing cross-time dependencies. When we now omit everything except for time mixing on Weather at 192, we suffer a 13.7% performance decrease. Additionally, substituting sLSTM blocks with mLSTM, LSTM, or GRU (#2–#4) consistently degrades performance, particularly for LSTM and GRU at longer horizons. We attribute this to the sLSTM’s inherent structure, which provides stronger mixing capabilities and mitigates the degradation commonly observed over extended prediction windows. However, we also observe that some configurations of xLSTM-Mixer, which exclude specific components, remain competitive. For instance, #7, which excludes the initial embedding token, still performs very well. This suggests that while it contributes positively to the overall performance, the model can sometimes still achieve competitive results without it. Similarly, depending on the dataset and target metric, initial forecasting with DLinear instead of NLinear is a viable option, too (#6).
The ablation study confirms that all components of xLSTM-Mixer contribute to its effectiveness, with the full configuration yielding the best results. Furthermore, we identified the sLSTM blocks and time-mixing components as critical for ensuring high accuracy across datasets and prediction lengths.